OpenStudy (calculusxy):

After driving 120 miles, a motorist found that he had covered 40% of his trip. How many miles did he have to complete his journey? A. 150 B. 300 C. 180 D. 20

OpenStudy (calculusxy):

@satellite73

OpenStudy (calculusxy):

@Lena772

OpenStudy (lena772):

B. 300 final answer

OpenStudy (calculusxy):

Can you explain why?

OpenStudy (calculusxy):

@Lena772

OpenStudy (lena772):

Sure

OpenStudy (anonymous):

the question can be interpreted as "120 is 40% of what number" translate directly in to math 120 is means \(=\) 40% means \(.4\) of means \(\times \) what number, pick a variable, i pick \(x\) \[120=.4\times x\]

OpenStudy (lena772):

120/40 = 3. 1% of his journey = 3 miles 100% of his journey =300 miles

OpenStudy (anonymous):

solve via \[120=.4x\iff x=120\div .4=1200\div4=300\]

OpenStudy (calculusxy):

Can you help me with another one?

OpenStudy (lena772):

Mine is a more simplified way, but if your teacher wants workings you should go with @satellite73

OpenStudy (calculusxy):

I'm fine with the easy way. LOL

OpenStudy (lena772):

:P

OpenStudy (calculusxy):

Can you help me with another one?

OpenStudy (lena772):

Yes just post it and tag me